OpenStudy (anonymous):

Based on the the equations of the parabolas, identify each parabola whose focus and vertex lie in different quadrants.

OpenStudy (anonymous):

y=-(x^2/8)+(x/4)+(23/8) x=-(y^2/36) - (5y/18) + (299/36) y=(x^2/32) + (x/4) - (13/2) x=(y^2/16) + (y/4) - (19/4) y=-(x^2/24) - (5x/12) + (95/24)

OpenStudy (anonymous):

x=-(y^2/16) - (y/4) +(11/4)

OpenStudy (anonymous):

.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@jane11509 ?

OpenStudy (jane11509):

Nope :/ sorry

OpenStudy (anonymous):

for the first one, \(y=-\dfrac{x^2}{8}+\dfrac{x}{4}+\dfrac{23}{8}\). From this, we can find the vertex of it, x\(= \dfrac{-b}{2a}= \dfrac{-(1/4)}{2(-1/8)}=1\) and then replace to get y=3. Hence the vertex is (1,3). The shape of the parabola is up-side down with vertex (1,3). we need do some more steps to figure out whether the focus and directrix are in the same quadrant or not. \(y=-\dfrac{1}{8}(x^2-2x-23)\\=-\dfrac{1}{8}(x^2-2x+1-1-23)\\=-\dfrac{1}{8}(x+2)^2+3\)

OpenStudy (anonymous):